Grade 12th passAlgebra

Find the measure of each exterior angle of a regular polygon of 9 sides

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Sum of angles a regular polygon having side n = (n-2)×180°
Sum of angles a regular polygon having side 9 = (9-2)×180°
= 7×180° = 1260°
Each interior angle=1260/9 = 140°
Each exterior angle = 180° - 140° = 40°
Or,
Each exterior angle = sum of exterior angles/Number of angles = 360/9 = 40°

Thanks